OpenStudy (anonymous):

10 to the what power is -3

OpenStudy (solomonzelman):

\[\huge\color{blue}{ 10^{-3}=\frac{1}{10^3} } \] just like anytime you have,\[\huge\color{blue}{ a^{-b}=\frac{1}{a^b} } \]

OpenStudy (solomonzelman):

and 10^3=1000 so what's your final answer?

OpenStudy (anonymous):

1/1000?

OpenStudy (solomonzelman):

Yes!!